OpenStudy (anonymous):

how to solve 1 --- 1+x ---- y

OpenStudy (anonymous):

you mean like this? \[ \Large{1\over{1+x\over y}} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \Large\frac{1\over1+x}{y\over1}={1\over 1+x}\times{1\over y} \]

OpenStudy (anonymous):

\[={1\over y(1+x)}\]

OpenStudy (anonymous):

thank you so much you are a life saver!

OpenStudy (anonymous):

you flip the denominator and multiply instead of divide

OpenStudy (anonymous):

yw